# S-EQA: Tackling Situational Queries in Embodied Question Answering

### v1

@@ Abstract
We present and tackle the problem of Embodied Question Answering (EQA) with Situational Queries (S-EQA) in a household environment. Unlike prior EQA work tackling simple queries that directly reference target objects and quantifiable properties pertaining them, EQA with situational queries (such as “Is the bathroom clean and dry?”) is more challenging, as the agent needs to figure out not just what the target objects pertaining to the query are, but also requires a consensus on their states to be answerable. Towards this objective, we first introduce a novel Prompt-Generate-Evaluate (PGE) scheme that wraps around an LLM’s output to create a dataset of unique situational queries, corresponding consensus object information, and predicted answers. PGE maintains uniqueness among the generated queries, using multiple forms of semantic similarity. We validate the generated dataset via a large scale user-study conducted on M-Turk, and introduce it as S-EQA, the first dataset tackling EQA with situational queries. Our user study establishes the authenticity of S-EQA with a high of the generated queries being deemed answerable, given the consensus object data. Conversely, we observe a low correlation of on the LLM-predicted answers to human-evaluated ones; indicating the LLM’s poor capability in directly answering situational queries, while establishing S-EQA’s usability in providing a human-validated consensus for an indirect solution. We evaluate S-EQA via Visual Question Answering (VQA) on VirtualHome, which unlike other simulators, contains several objects with modifiable states that also visually appear different upon modification — enabling us to set a quantitative benchmark for S-EQA. To the best of our knowledge, this is the first work to introduce EQA with situational queries, and also the first to use a generative approach for query creation. Our work introduces a new task and approach that aims to improve the usability of embodied agents in real-world household environments.
@@ I Introduction
In contrast, answering queries such as “Is the kitchen ready for meal preparation?” requires a sense of situational awareness in gauging the status of multiple objects in the environment. While not uniformly defined, a consensus does exist on standard conditions that can be used to answer such questions. This refers to a set of generally accepted object states and conditions (such as a Knife being Present) for successfully evaluating the query. A household robot deployed in the real world is likely to encounter such queries that require it to utilize consensus knowledge pertaining which objects and associated states are commonly associated with the situation. We term these as situational queries, and in this work, tackle the twofold challenge of 1) conjuring authentic situational queries and corresponding consensus object data and 2) highlighting the role of the consensus in simplifying the process of responding to situational queries.
@@
A query such as “Is the house ready for sleeptime?” is a situational query, since there exist generally accepted object states, such as dimmed lights, closed windows, locked doors, etc., which most people associate with a house being ready for nighttime. However, a query like “Can I bring guests over?”, while situational, is too ambiguous, as the answer relies on various unstated factors and user preferences on household norms, and event-specific constraints. For situational queries to be validatable, it is important for them to not be too ambiguous so as to allow for a consensus to exist. Our work places emphasis on this, and follows an in-context learning scheme on an LLM to avoid overly simplistic nor excessively ambiguous queries.
@@
Developing a dataset for EQA in the past has been limited to constructing simple queries using entities from a simulator [4, 5], with some work extending it to abstract queries built on scene-graphs [6]. Posing situational queries using these existing techniques is however nontrivial, as they stem from human intuition and contextual awareness of the environment. Setting up real-world experiments to collect this data is also cumbersome; if not having an intrusive agent capturing conversations at home, an offline approach would be cognitively heavy, requiring humans to come up with queries for imaginary situations. As such, in our work, we use an LLM to generate and study such queries, while ensuring they cover unique concepts stemming from various situations. Table I compares our dataset with existing ones.
Main Results: To tackle these issues, we leverage the commonsense capabilities of Large Language Models (LLMs) and present an approach to generate situational queries and associated consensus object states and relationships. LLMs allow us to generate situational query data for a variety of scenarios which are nontrivial to obtain via real-world data collection. The main contributions of our work are as follows:-
We propose a novel Large Language Model (LLM)-based method for generating Situational Queries for Embodied Question Answering (EQA). Unlike simple queries that directly reference particular objects, situational queries necessitate an understanding of the consensus objects and states pertaining a situation.
Our proposed Prompt-Generate-Evaluate (PGE) scheme wraps around an LLM to generate a wide array of unique and challenging queries covering various household scenarios. Along with the queries, we also generate corresponding consensus object information, pointing to object states and relationships relevant to the queries.
The generated queries are validated via a large scale MTurk [7] user study to produce S-EQA, a LLM-generated dataset containing situational queries, associated consensus object data, and predicted answers. While the number of queries is small due to their highly abstract nature, PGE ensures diversity in the queries across rooms.
We quantitatively evaluate S-EQA by performing Visual Question Answering (VQA) on VirtualHome images with both situational queries (Room VQA) and consensus data (Object VQA) to establish a benchmark.
We observe that 97.26% of users find the generated queries to be answerable provided the consensus data, while only 46.2% of user answers align with the LLM predicted ones. From the former, we can infer the authenticity of LLM generated data in producing human-like consensus states, while the latter indicates the LLM cannot be used to directly answer situational queries; establishing the need and use-case for S-EQA, a human-validated dataset that provides situational queries along with consensus objects and states.
@@ II Related Work
Sintean et. al in K-EQA [6] attempt to tackle this problem, by using scene-graph data and formal grammar to postulate queries. However, these queries are still limited to simple object grounding within a scene, and do not require any form of situational awareness about the readiness of an environment. For instance, S-EQA considers queries like “Is the living room ready for a movie night?”, and presents consensus multiple object states (TV: On, Popcorn: Cooked, On Table, etc.) to measure this readiness. For the same example, a K-EQA query would have been constructed as “Is there a device for watching movies in the living room?”, which requires only scene knowledge for an answer, but does not require situational knowledge about the consensus states of objects for a movie night.
In our work, we utilize an LLM to generate situational queries and approximate human consensus on object states required to answer the query. Our queries are an abstraction above the scene-knowledge based K-EQA queries, requiring consensus knowledge about multiple object states for validation.
LLMs for Query Generation: A recent trend has seen the emergence of Large Language Models (LLMs) such as GPT-3.5 [11], GPT-4 [12], LLaMA [13], Palm [14] and Palm-e [15]. These models are well known for their commonsense reasoning capabilities [16], especially in the context of question-answering. In embodied AI, these capabilities have recently been leveraged to produce superior planning and decision-making strategies during navigation [17, 18, 19, 20]. In our case, we wish to generate situational queries for embodied question answering, which as defined in Section I require consensus about object states and relationships in the environment. As such, we are motivated to use LLMs for generating these queries and ask if their commonsense reasoning capabilities can sufficiently represent the consensus needed to answer situational queries.
III Situational EQA: Generation, Analysis, and Evaluation
Having motivated the problem of Situational EQA, we now present a detailed description of our approach. We present three phases — generating situational data using LLMs, validating this data via segregation and a user survey, and finally evaluating it on the VirtualHome simulator. We hope the generative approach being used to develop S-EQA helps foster future research in this domain.
@@ III-A Situational Query Definition
Situational queries in our work are defined as questions requiring the assessment of a collection of consensus objects and states to reach an existential ‘Yes/No’ answer. For example, a query such as “Is the livingroom ready for movie night?” has a Yes/No answer based on whether consensus objects like the TV, lights, etc. are in their appropriate state. Additionally, we want to ensure that the queries are related to a situation and are not simple in nature, i.e., the target objects and states are not in the query itself. For example, “Is the sofa blue?” [1] or “How many objects can be used to hold rice?” [6] are simple queries, as they do not deal with a situation, but rather pose a query referencing objects in the environment. In contrast, “Is the dining area set up for dinner?” is situational, as it doesn’t explicitly mention any objects, but requires the identification of relevant objects (eg. table, napkins), and a consensus on what their states should be (open, present) for the answer. Considering the LLM as a function that generates queries , and to be a set of all possible queries generated by , we want situational queries , which satisfies the following conditions -
The query must not directly reference any object, i.e., it must omit words from a set which includes “object”, “table”, “chair”, etc.
The query must be a question.
The answer to the query must be binary i.e., Yes/No, and not subjective in nature.
@@ III-B Generating Situational Queries: S-EQA
Figure 2 presents an overview of our generation approach. We generate datapoints of the format - [Situational Query, Object-States, Object-Relationships, LLM Answer]. A sample datapoint is as follows -
Query: “Is the bathroom ready for a shower?”
@@
LLM Answer: ‘Yes’
Here, the LLM points to a set of object states and relationships (such as ‘towels inside bathroom being present’) that it believes need to be checked to determine if the bathroom is ready for a shower. We also predict an answer to the query that is later verified in our user study.
III-B1 VirtualHome Setup
We use the VirtualHome [21] simulator for defining ground truth objects, states and properties in our experiments. It contains 7 different household environments containing over 390 objects. Unlike other embodied simulators, VirtualHome is unique in that it contains several household objects including interactable ones, with easily modifiable locations and states. The objects are also visually different upon modification, and this allows us to perform VQA later on to evaluate our dataset.
@@ III-B1 VirtualHome Setup
III-B2 Prompt-Generate-Evaluate (PGE) Scheme
Figure 2 presents an overview of our Prompt-Generate-Evaluate scheme for generating data. We perform analytical operations on the LLM’s generated datapoints to ensure they fit into our definition of situational queries, while also maintaining diversity. The System Prompt for our LLM contains a list of all object states and relationships. We also provide additional rules and in-context examples to improve the generation of prompts. For instance, explicitly mentioning “Do not conjure any new object states” ensures that the generated datapoints do not contain states that are not achievable in VirtualHome. Additionally, we also provide a set of “Generated Queries” as feedback to the system prompt. These represent queries that the LLM has already generated, and in the User prompt, encourage the LLM to generate queries different from these. Our strategy for choosing these queries is discussed in the feedback measures presented below.
Our User Prompt asks to generate datapoints in the format specified earlier, and provides a few sample datapoints. We also provide some in-context rules to ensure that the generated queries are not too simplistic, not too ambiguous, and are also unique. For instance, we explicitly state ‘The situational query must not reference any object or even contain the word ‘object’.’ and ‘The query must have a Yes/No answer.’. Additionally, we ask it to generate queries very different from what is already present in the Generated Queries in the System Prompt, to avoid simple sentence rephrasing.
To ensure that our Situational Query Database contains data points that adhere to the specifications presented above, we take the following two feedback measures:
After generating data points, we embed the situational queries using RoBERTa [22] to get sentence embeddings . We perform agglomerative clustering [23] on to accumulate them into groups of queries based on sentence similarity. We then pick the query closest to the centroid of each of these clusters to get queries that best represent the diversity in the generated data. Finally, we update the system prompt with these queries.
Simultaneously, we compute the cosine similarity of the generated query embeddings with the queries already present in the Situational Query Database . If is greater than a particular threshold (), we continue the “conversation” with the LLM, by presenting it with a re-generation prompt. We mention the percentage of similarity, and ask the LLM to re-generate a new batch of datapoints.
These feedback measures ensure contains unique situational queries that are not simple rewordings, while also encouraging diversity in the generated situations.
@@ III-B2 Prompt-Generate-Evaluate (PGE) Scheme
III-C Validating S-EQA: t-SNE Plots and MTurk Survey
Our objective with validating S-EQA is not just to measure the authenticity of the generated data, but also to establish necessity and the real-world usability of our dataset. First, we first visualize the queries in using t-SNE plots to analyze their diversity. We then conduct a Mechanical Turk based user study to validate the generated data.
III-C1 Analyzing S-EQA with t-SNE
Figure 3 illustrates t-SNE plots we derive from . These are as follows -
Room Categorization: Using the room properties obtained from VirtualHome (section III-B1), we group the questions into 6 categories - [kitchen, livingroom, bedroom, bathroom, multi-room, no room] (Ref. Fig. 3(a)). We infer from this plot that most questions being generated belong to one particular room, as opposed to multiple ones. This suggests that the LLM tends to generate situational queries that are usually room-specific, referring to multiple objects in a single area. As such, multi-room scenarios (like “Is the house ready for sleeptime?”) are less common in the generated data.
Situational Categorization: Here, we use the LLM as a classifier to predict if the generated questions are situational according to our definition in III-A. We prompt and present in-context examples accordingly. We notice that of generated questions (Ref. Fig. 3(b)) are classified as situational according to the LLM. As the LLM classifications have several false positives, this result motivates us to use human-validation to authenticate the dataset.

### v3

# Is the House Ready For Sleeptime? Generating and Evaluating Situational Queries for Embodied Question Answering

@@ Abstract
We present and tackle the problem of Embodied Question Answering (EQA) with Situational Queries (S-EQA) in a household environment. Unlike prior EQA work tackling simple queries that directly reference target objects and properties (“What is the color of the car?”), situational queries (such as “Is the house ready for sleeptime?”) are challenging as they require the agent to correctly identify multiple object-states (Doors: Closed, Lights: Off, etc.) and reach a consensus on their states for an answer. Towards this objective, we first introduce a novel Prompt-Generate-Evaluate (PGE) scheme that wraps around an LLM’s output to generate unique situational queries and corresponding consensus object information. PGE is used to generate datapoints in the VirtualHome simulator, which is then annotated for ground truth answers via a large scale user-study conducted on M-Turk. With a high rate of answerability () on this study, we establish that LLMs are good at generating situational data. However, in evaluating the data using an LLM, we observe a low correlation of with the ground truth human annotations; indicating that while LLMs are good at generating situational data, they struggle to answer them according to consensus. When asked for reasoning, we observe the LLM often goes against commonsense in justifying its answer. Finally, we utilize PGE to generate situational data in a real-world environment, exposing LLM hallucination in generating reliable object-states when a structured scene graph is unavailable. To the best of our knowledge, this is the first work to introduce EQA in the context of situational queries and also the first to present a generative approach for query creation. We aim to foster research on improving the real-world usability of embodied agents through this work.
@@ I Introduction
In contrast, answering queries such as “Is the kitchen ready for meal preparation?” requires a sense of situational awareness in gauging the status of multiple objects in the environment. While not uniformly defined, a consensus does exist on standard conditions that can be used to answer such questions. This refers to a set of generally accepted object states and conditions (such as a Knife being Present) for successfully evaluating the query. A household robot deployed in the real world is likely to encounter such queries that require it to utilize consensus knowledge pertaining which objects and associated states are commonly associated with the situation. We term these as situational queries, and in this work, tackle the twofold challenge of 1) conjuring authentic situational queries and corresponding consensus object data and 2) highlighting the role of consensus object states in simplifying the process of responding to situational queries.
@@
A query such as “Is the house ready for sleeptime?” is a situational query, since there exist generally accepted object states, such as dimmed lights, closed windows, locked doors, etc., which most people associate with a house being ready for nighttime. However, a query like “Can I bring guests over?”, while situational, is too personal, as the answer relies on various user preferences on household norms and event-specific constraints. In our work, we aim to generate and human-validate queries where a general consensus exists on multiple object states.
@@
Developing EQA datasets in the past has been limited to constructing simple queries using entities from a simulator [5, 6], with some work extending it to abstract queries built on scene-graphs [7]. Posing situational queries using these existing techniques is however nontrivial, as they stem from human intuition and contextual awareness of the environment. Setting up real-world experiments to collect this data is also cumbersome; if not having an intrusive agent capturing conversations at home, an offline approach would be cognitively heavy, requiring humans to imagine situations for queries. In our work, we develop a scheme that utilizes a Large Language Model (LLM) to generate such queries, providing in-context cures to cover unique concepts stemming from various situations. Table I compares our dataset with existing work.
Main Results: We tackle the issues presented above, utilizing LLMs as a prior for generating situational queries and associated consensus object states and relationships111https://youtu.be/m3hg59zGYg8. We human-validate the generated data, and present a benchmark for EQA on situational queries. The main contributions of our work are as follows:-
Problem: We tackle the novel problem of posing and evaluating Situational Queries for Embodied Question Answering (S-EQA). Unlike simple queries that directly reference particular objects, posing situational queries itself is challenging, requiring a consensus on multiple objects and their states, and cannot be formed easily using prior rule-based methods (Refer Table I).
Method: We propose a novel Prompt-Generate-Evaluate (PGE) scheme that integrates an LLM with a feedback scheme that allows for the generation of several unique situational queries, along with corresponding consensus object-state and object-relationship information that needs to be satisfied to successfully answer the question.
Human-Validation: Using PGE in the VirtualHome simulator, we produce a S-EQA dataset of situational queries and consensus data. We annotate these via a large scale MTurk [8] user study, which gives us a high authenticity percentage of , indicating that the data generated corresponds to consensus.
LLM Eval: We prompt the LLM to provide binary answers to S-EQA datapoints generated from VirtualHome. Despite being able to generate highly authentic situational data, we observe a subpar correlation of with the human annotations when it comes to answering. We infer that LLMs can generate good situational queries and consensus object-states, but answer them poorly.
LLM Reasoning: When asked for an explanation, we observe the poor commonsense reasoning of LLMs in justifying their answer. This further suggests that LLMs exhibit poor situational awareness.
Real World Challenges: Finally, we discuss the transfer of PGE to a real world real world lab environment, and discuss various challenges with LLM hallucination that obstruct the generation of quality consensus data.
@@ II Related Work
Sintean et. al in K-EQA [7] attempt to tackle this problem, by using scene-graph data and formal grammar to postulate queries. However, these queries are still limited to simple object grounding within a scene, and do not require any form of situational awareness about the readiness of an environment. For instance, S-EQA considers queries like “Is the living room ready for a movie night?”, that requires consensus on multiple objects and their states (TV: On, Popcorn: Cooked, On Table, etc.) to measure this readiness. A K-EQA query for the same example could be constructed as “Is there a device for watching movies in the living room?”, which requires only scene knowledge about the availability of a TV, but does not require situational knowledge about the various object-states involved in a movie night.
In our work, we utilize an LLM to generate situational queries and approximate human consensus on object states. Our queries are an abstraction above the scene-knowledge based K-EQA queries, requiring consensus knowledge about multiple object states for validation.
LLMs for Query Generation: A recent trend has seen the emergence of Large Language Models (LLMs) such as GPT-3.5 [12], GPT-4 [13], LLaMA [14], Palm [15] and Palm-e [16]. These models are well known for their commonsense reasoning capabilities [17], especially in the context of question-answering. In embodied AI, these capabilities have recently been leveraged to produce superior planning and decision-making strategies during navigation [18, 19, 20, 21]. In our case, we wish to generate situational queries for embodied question answering and ask if the common sense reasoning capabilities of LLMs can both generate such queries as well as the the consensus object-states needed for sufficiently answering them.
III Situational EQA: Generation & Validation
Having motivated the problem of Situational EQA, we now present a detailed description of our approach. We present three phases — generating situational data using an LLM (GPT-4), annotating the data via a large-scale user survey to produce a S-EQA dataset, and finally evaluating it on the VirtualHome simulator.
@@ III-A Situational Query Definition
Situational queries in our work are defined as questions requiring the assessment of a collection of consensus objects and states to reach an existential ‘Yes/No’ answer. For example, a query such as “Is the livingroom ready for movie night?” has a Yes/No answer based on whether consensus objects like the TV, lights, etc. are in their appropriate state. Additionally, we want to ensure that the queries are related to a situation and are not simple in nature, i.e., the target objects and states are not referenced in the query itself. For example, “Is the sofa blue?” [2] or “How many objects can be used to hold rice?” [7] are simple queries, as they do not deal with a situation, but rather pose a query referencing objects in the environment. In contrast, “Is the dining area set up for dinner?” is situational, as it doesn’t explicitly mention any objects, but requires the identification of relevant objects (eg. table, napkins), and a consensus on what their states should be (open, present) for the answer. Considering the LLM as a function that generates queries , and to be a set of all possible queries generated by , we want situational queries , which satisfies the following conditions -
Abstraction: The query must not directly reference any object, i.e., it must omit words from a set containing words such as “object”, “table”, “chair”, etc.
Contextual: The query requires an assessment of relevant objects and their states collectively satisfy a scenario (e.g., “Is the dining area set up for dinner?”), requiring identification of appropriate objects and consensus on their expected states (e.g., table present, napkins, lighting, etc.).
Binary: The answer to the query must be binary i.e., Yes/No, and not subjective in nature.
@@ III-B Generating Situational Queries: S-EQA
We prompt the LLM to generate datapoints of the format - [Situational Query, Object-States, Object-Relationships]. For example,
Situational Query:
“Is the bathroom ready for a shower?”
@@
The object states and relationships (such as ‘towels inside bathroom being present’) being generated represent scene-graph requirements that the LLM believes need to be checked for the query to satisfy equation 1.
VirtualHome Setup: We use the VirtualHome [1] simulator for defining ground truth objects, states and properties in our experiments. It contains 7 different household environments containing over 390 objects. Unlike other embodied simulators, VirtualHome is unique in that it contains several household objects including interactable ones, with easily modifiable locations and states. The objects are also visually different upon modification, and this allows us to perform VQA later on to evaluate our dataset.
@@
Prompt-Generate-Evaluate (PGE) Scheme: Figure 2 presents an overview of our Prompt-Generate-Evaluate scheme for generating data. We analyze the LLM’s generated datapoints to provide feedback for ensuring the data fits into our definition of situational queries, while also maintaining diversity.
The System Prompt for our LLM contains a list of all object states and relationships, which in this case is the scene-graph provided by VirtualHome. We also provide additional rules and in-context examples to improve the generation of prompts. For instance, explicitly mentioning “Do not conjure any new objects or states” ensures that the generated data does not contain objects not present or states that are not achievable in VirtualHome. Additionally, we also provide a set of “Generated Queries” as feedback to the system prompt. These queries are representative of what the LLM has already generated, and we encourage the LLM to generate queries different from these.
Our User Prompt asks to generate datapoints in the format specified earlier, and while providing a few manually annotated sample datapoints. Additionally, we provide some in-context rules to ensure that the generated queries are not too simplistic, not too ambiguous, and are also unique. For instance, we explicitly state ‘The situational query must not reference any object or even contain the word ‘object’.’ and ‘The query must have a Yes/No answer.’. We also ask it to generate queries very different from what is already present in the Generated Queries in the System Prompt and come up with ‘unique and creative household situations‘ with the object data provided.
To ensure that our Situational Query Database contains diverse data points adhering to specifications presented in equation 1, we take the following two feedback measures:
Generated Query Update: After generating data points, we embed the situational queries using RoBERTa [22] to get sentence embeddings . We perform agglomerative clustering [23] on to accumulate them into groups of queries based on sentence similarity. We then pick the query closest to the centroid of each of these clusters to get queries that best represent the diversity in the generated data. Finally, we update the system prompt with these queries.
Continued Conversation: Simultaneously, we compute the cosine similarity of the generated query embeddings with the queries already present in the Situational Query Database . If is greater than a particular threshold (), we continue the conversation with the LLM, by presenting it with a re-generation prompt. We mention the percentage of similarity, and ask the LLM to re-generate a new batch of datapoints.
These feedback measures ensure contains unique situational queries that are not simple rewordings, while also encouraging diversity in the generated situations. Contrasting examples with and without PGE are illustrated in Fig. 4.
@@
Prompts: We use the following prompts for PGE.
System Prompt:
“I have a list of objects, and their states, and relationships in a household.
Object states are listed in the OBJ_STATE_DICT dictionary below. Each item has the format OBJECT: [STATES].
STATES, IF PRESENT, can switch between ON/OFF and OPEN/CLOSED. Do not conjure any nrew objects or states.
OBJ_REL_DICT contains the initial relationships between the objects. These can be changed. For instance, "apple INSIDE fridge" is a valid relation.
OBJ_STATE_DICT :- {…}
OBJ_REL_DICT :- {…}
GENERATED_QUERIES :- {…} ”
User Prompt:
“ Using OBJ_STATE_DICT and OBJ_REL_DICT, can you generate {X} potential questions, states and relationships that the user might ask about the environment?
These must be about a potential scenario, requiring situational awareness and a consensus on multiple object, their states and relationships.
The output should be lines of the form [Question, Object-State Pairs, Relationships].
Make sure you generate questions very different from those in GENERATED_QUERIES. Get creative with the potential scenarios!
Make sure to use an exhaustive set of relationships and object-state pairs for each query.
The query must have a Yes/No answer.
The query must not directly reference any object or even contain the word ‘object’. ”
Re-Generation Prompt:
“{X}% of the questions are similar to what you’ve already generated earlier! Try again, give me {Y} more.
QUERIES:…”
III-C Validating S-EQA: t-SNE plots and M-Turk Annotation
Our objective with validating S-EQA is not just to annotate the queries, but to also measure the authenticity of the generated data. We first visualize the queries in using t-SNE plots to analyze their diversity. We then conduct a user study via Mechanical Turk [8] to annotate the generated data.
t-SNE Plots: We embed the generated queries with BERT and visualize them using t-SNE. They are then are categorized by Room (pick from [kitchen, living room, bedroom, bathroom, multi-room, no room]), Situational/Simple (Yes/No) and Spatio-Temporal (time-sensitivity). We observe that GPT-4 tends to generate queries specific to objects in a single room over multi-room or no-room ones like “Is the house ready for sleeptime?”. Further, we note that most queries () are classified as situational and not simple. Around are categorized as spatial queries (like “Is the kitchen ready for cooking?”) that refer to the current environment state, as opposed queries like “Was someone in the kitchen today?” which require deeper situational understanding involving time. Overall, we observe that most queries adhere to the definition provided in Equation 1.
Room Categorization: Using the room properties obtained from VirtualHome (section III-B), we group the questions into 6 categories - [kitchen, livingroom, bedroom, bathroom, multi-room, no room]. We infer from this plot that most questions being generated belong to one particular room, as opposed to multiple ones. This suggests that the LLM tends to generate situational queries that are usually room-specific, referring to multiple objects in a single area. As such, multi-room scenarios (like “Is the house ready for sleeptime?”) are less common in the generated data.
Situational Categorization: Here, we use the LLM as a classifier to predict if the generated questions are situational according to our definition in III-A. We prompt and present in-context examples accordingly. We notice that of generated questions are classified as situational according to the LLM. As the LLM classifications have several false positives, this result motivates us to use human-validation to authenticate the dataset.